OpenStudy (anonymous):

Sarah's craft project uses pieces of yarn that are 1/18 yard long. she has a piece of yarn that is 3 yards long .how many 1/8 yards pieces can she cut and still have 1 1/4 yards left?

OpenStudy (phi):

to find how many times 1/8 "goes into" 3 you divide 1/8 into 3 in other words you do \[ \frac{3}{\frac{1}{8} }\] one way to do that is multiply top and bottom by 8 \[ \frac{3\cdot \frac{8}{1}}{\frac{1}{8} \cdot \frac{8}{1}}\] can you simplify that ?